OpenStudy (xxuan):

1. A right circular cylinder has a base diameter and height of 10 inches each. What is the total surface area of this cylinder, in square inches? a. 25π b. 50π c. 75π d. 150π e. 200π 2. The interior dimensions of a rectangular box are 5 inches by 4 inches by 3 inches. What is the volume, in cubic inches, of the interior of the box? a. 12 b. 60 c. 90 d. 120 e. 124

TheSmartOne (thesmartone):

\(\bf\Large SA=2\pi h+2\pi r^2\)

TheSmartOne (thesmartone):

Use that for the first question.

TheSmartOne (thesmartone):

Use this for the second question: \(\bf\Large V = Length\times Height \times Width\) Basically, you multiply all the three numbers they gave you together to get your answer.

OpenStudy (xxuan):

number 1 radius is 5 ?

Directrix (directrix):

For the first problem, the surface area formula should be this: SA = 2*π*r*h + 2* π* r²

Directrix (directrix):

On that problem, r = 5 and h = 10 SA = 2*π*5*10 + 2* π* 5² SA = ? @XXUAN

Directrix (directrix):

>>number 1 radius is 5 ? Yes.

TheSmartOne (thesmartone):

My bad, I accidentally removed that r in the equation. Yes, the radius is 5. :)